In a series R-L-C circuit the voltage amplitude across the resistor is 35.0 V, across the capacitor is 95.0 V and across the inductor is 55.0 V. When the source sinusoidal frequency is 10.6 KHz and the resistor value is 60.0 ohms

 A. What is the capacitance of the capacitor?
Answer #1

the voltage across each component can be described mathematically according to the current flowing through, and the voltage across each element as.

Instantaneous Voltages for a series RLC Circuit

From the voltage across the resistance we obtain the magnitude of the current:

35.0V=I60.0\Omega\rightarrow I=0.58A

It's a series circuit so current is the same through the whole circuit, then:

95.0V=\frac{0.58A}{(2\pi10.6\cdot 10^3Hz)C}\rightarrow C=9.17\cdot 10^{-8}F\approx 0.092 \mu F
Answer #1

the voltage across each component can be described mathematically according to the current flowing through, and the voltage across each element as.

Instantaneous Voltages for a series RLC Circuit

From the voltage across the resistance we obtain the magnitude of the current:

35.0V=I60.0\Omega\rightarrow I=0.58A

It's a series circuit so current is the same through the whole circuit, then:

95.0V=\frac{0.58A}{(2\pi10.6\cdot 10^3Hz)C}\rightarrow C=9.17\cdot 10^{-8}F\approx 0.092 \mu F